OpenStudy (anonymous):

verify this trig identity and scan in if possible plz... need all help i can get... :)

OpenStudy (anonymous):

\[(tanxcotx/cosx)=secx \]

OpenStudy (anonymous):

tan x=sin x/cos x and cot x = 1/tan x= cos x/sin x so, (tan x)(cot x)=(sin x)(cos x)/(cos x)(sin x)= 1 therefore i/cos x= sec x

OpenStudy (anonymous):

Here's another approach. Since we know tanxcotx = 1. \[\left( tanxcotx\right)/cosx =secx\] \[1/cosx = secx \]